OpenStudy (anonymous):

(5/t-2) - (3t/t-2) = 4/t^2-4t+4

OpenStudy (lalaly):

\[(\frac{5}{t-2} \times \frac{t-2}{t-2})-(\frac{3}{t-2} \times \frac{t-2}{t-2}) =\frac{4}{(t-2)(t-2)}\]\[\frac{5t-10-3t+6}{(t-2)(t-2)}=\frac{4}{(t-2)(t-2)}\]\[2t-4=4\]\[2t=8\]

OpenStudy (lalaly):

t=4

OpenStudy (anonymous):

isnt it going to be 3t^2? It's 3t(t-2)

OpenStudy (lalaly):

oops my bad

OpenStudy (lalaly):

\[\frac{5t-10-3t^2+6t}{\cancel{(t-2)(t-2)}}=\frac{4}{\cancel{(t-2)(t-2)}}\]

OpenStudy (lalaly):

\[-3t^2+11t-10=4\]

OpenStudy (lalaly):

\[-3t^2+11t-14=0\]use the quadratic formula now

OpenStudy (anonymous):

uhhh i got \[\left(\begin{matrix}-11\pm \sqrt{-47} \ \\ -6\end{matrix}\right)\]